OpenStudy (anonymous):

evaluate the integral: sin^2 (k) dx between 0 and 2pi

OpenStudy (anonymous):

write it as: \[sin^2(x) = \frac{1-cos2x}{2}\]

OpenStudy (anonymous):

sorry its actually : sin^2 (k)x dx between 0 and 2pi

OpenStudy (anonymous):

If \(k\) is a constant then treat \(sin^2(k)\) just like you would any other number because it will also be constant. So you will get: \[\frac{sin^{2}(k)x^2}{2}\] between 0 and \(2\pi\). Are you able to finish it off?

OpenStudy (anonymous):

Also do you see how we get that? It's just like integrating \(3x\), which would be \(\frac{3x^2}{2}\).

OpenStudy (anonymous):

wait how did you get x^2

OpenStudy (anonymous):

Would you be able to integrate something like \(2x\)?

OpenStudy (anonymous):

yes i can

OpenStudy (anonymous):

Is \(k\) just a constant in your question btw I'm assuming?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

If \(k\) is just a constant then \(sin^2(k)\) will also be a constant, so just treat \(sin^2(k)x\) like you would any other constant multiplying \(x\). Does that make sense?

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

Anything I can help with?

OpenStudy (anonymous):

nope. i got it. thank you :D

OpenStudy (anonymous):

Excellent :)